OpenStudy (anonymous):

State whether the slope of the line passing through (2, 4) and (5, 2) is positive, negative, zero, or undefined.

OpenStudy (anonymous):

negative

OpenStudy (anonymous):

To find any slope we use the formula y2-y1/x2-x1. Pluck in the numbers into this and solve. 2-4/5-2=-2/3 -2/3 is our slope so that means it is negative

jimthompson5910 (jim_thompson5910):

Slope m = (y2-y1)/(x2-x1) m = ( 2- 4)/(5-2) m = -2/3 So the slope of the line that passes through the two points is -2/3, which is negative